OpenStudy (anonymous):

Cadence spent $42.40 on short-sleeve and long-sleeve T-shirts. The long-sleeve T-shirts cost $14 each. If she bought two of each type of shirt, what was the price of each short-sleeve T-shirt?

jagr2713 (jagr2713):

is there an answer choice

jagr2713 (jagr2713):

(14x2 - 28) (28- 42.40 =14.40 divided by 2 = 7.20)

OpenStudy (anonymous):

here are the chocies

OpenStudy (anonymous):

a.7.00 b.7.10 c.7.15 d.7.20

jagr2713 (jagr2713):

So its D

jagr2713 (jagr2713):

(14x2 - 28) (28- 42.40 =14.40 divided by 2 = 7.20)

jagr2713 (jagr2713):

you get it

OpenStudy (anonymous):

yea and thank you so much for your help

jagr2713 (jagr2713):

Yw

OpenStudy (texaschic101):

2(14) + 2s = 42.40 (s stands for short sleeve shirts) 28 + 2s = 42.40 2s = 42.40 - 28 2s = 14.40 s = 14.40/2 s = 7.20